OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients. y = -6.9x + 5.1

OpenStudy (anonymous):

A. 69x + 10y = 51 B. -69x + 10y = -51 C. 69x + 10y = -51 D. -69x + 10y = 51

OpenStudy (anonymous):

1. multiply both sides by 10. 2. make the equation into Ax+By = C

OpenStudy (anonymous):

did you figure it out?

OpenStudy (anonymous):

nope im trying to do it now. Algebra 2 is not my strongest

OpenStudy (anonymous):

I will help you. just follow the steps that I said. did you multiply it?

OpenStudy (anonymous):

lol. do i multiply the -6.9x times 10 and the y times 10

OpenStudy (anonymous):

im lost now

OpenStudy (anonymous):

Sorry I changed the numbers awhile ago: \[(10 \times (y)) =( -6.9x + 5.1)\times10\] after you multiply it by 10, the equation will be: \[10y = -69x + 51\]

OpenStudy (anonymous):

ohokay thats how i was about to solve it thank you

OpenStudy (anonymous):

REMEMBER: When you transpose(or move a) term on the other side of the equation, the sign of that term will change. What's your answer?

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Nope... you move -69x on the other side

OpenStudy (anonymous):

so it will change its sign to positive

OpenStudy (anonymous):

C

OpenStudy (anonymous):

NOpe...don't guess

OpenStudy (anonymous):

51 was not transposed

OpenStudy (anonymous):

Why is that

OpenStudy (anonymous):

because you want Ax+By =C The equation before is in the form y=mx + b or By = Ax +C make it Ax+By =C Ax = 69x By = 10y C= 51

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

@andrealane92

OpenStudy (anonymous):

Yes thanks

OpenStudy (anonymous):

what is your answer?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

yup